OpenStudy (anonymous):

solve the second compound inequality. graph the solution. -2 less than or equal 2x -14 < 4

OpenStudy (anonymous):

I know the answer is 1 less than or equal to x <4 but i need the work part.

OpenStudy (anonymous):

i got 6 less than or equal to x <9

OpenStudy (anonymous):

\[6 \le x < 9\]

OpenStudy (anonymous):

whoops i meant -2 less than or equal to 2x-4 <4

OpenStudy (anonymous):

thats the real question

OpenStudy (anonymous):

\[1 \le x < 4\]

OpenStudy (anonymous):

use the same example as above

OpenStudy (anonymous):

@nadialabish94 can you help? its not letting me open his attachment

OpenStudy (anonymous):

sure. the first thing you do is separate it into two problems so... \[-2 \le 2x-4\] and \[2x-4 < 4\] Then you solve for x.

OpenStudy (anonymous):

let me know what you get...

OpenStudy (anonymous):

1 <x < 4

OpenStudy (anonymous):

make sure that you keep the first sign \[\le \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome :)